OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. A. y = x + 2 B. y = -x - 2 C. y = -x + 4 D. y = x - 4

OpenStudy (anonymous):

help

OpenStudy (anonymous):

@SolomonZelman @mohammad.arqum

OpenStudy (anonymous):

y=-x+c 1=-x-3 x=-4 y=-x-4 i dont see this in the options i cant be wrong now can I :O

OpenStudy (anonymous):

So D ?

OpenStudy (solomonzelman):

Okay, so when we are talking about a parallel line going through a certain point, we are not changing the slope only, we are only changing the y-intercept it is going through. (In other words, we are shifting the line vertically up or down). I'll assume you know what parallel lines look like. And if you change the slope, you will change their slope, (or their incline) and this will make them intercept). So, we will need to solve the equation for the y-intercept. \(\large\color{black}{ y = -x + b. }\) It has to be going through \(\large\color{black}{ (-3,1) }\), So plug it ito the equation and solve for b.

OpenStudy (anonymous):

@SolomonZelman man try to put the values into the equation im not getting a correct answer.!

OpenStudy (anonymous):

So B or D

OpenStudy (anonymous):

I think it's b

OpenStudy (solomonzelman):

I'll plug it in for you, and you do it, @IndianBoyNiko okay? \(\large\color{black}{ (1)=-(-3)+b}\) can you solve for b? I am sure B is correct though, but you seemed to just guess it...

OpenStudy (anonymous):

I did the test before, i was just seeing if i got the answer right

OpenStudy (solomonzelman):

B is right.

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

can i ask another

OpenStudy (anonymous):

OOOHHH silly mistake sorry about that.... y=-x+c 1=-(-3)+c c=-2 y=-x-2 so its B

OpenStudy (anonymous):

i'm just gonna ask it

OpenStudy (anonymous):

If f(x) is the total cost, in dollars, of x candies, which of the following statements best describes the meaning of f(2) = 6? A. The total cost of 2 candies is $6. B. The total cost of 6 candies is $2. C. The total cost of 2 candies is $3. D. The total cost of 3 candies is $2.

OpenStudy (anonymous):

Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) - 3? A. The graph shifts 3 units up. B. The graph shifts 3 units down. C. The graph shifts 3 units left. D. The graph shifts 3 units right.

OpenStudy (anonymous):

@IndianBoyNiko Mass tagging is against the Code of Conduct,

OpenStudy (anonymous):

holy cow thats alot of people lol

OpenStudy (anonymous):

Oh sorry

OpenStudy (anonymous):

Who cares. He obviously needs help!

OpenStudy (anonymous):

use mathway.com

OpenStudy (anonymous):

I just need two questions answered so i can see if i'm right

OpenStudy (anonymous):

Don't mass tag

OpenStudy (anonymous):

help ?

OpenStudy (anonymous):

i think that answer is 'B' @IndianBoyNiko

OpenStudy (anonymous):

For which question ? @shereenkhan

OpenStudy (anonymous):

Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) - 3? this question brother :)

OpenStudy (solomonzelman):

you want ti tnaswered, or explained? I'll do the second only:)

OpenStudy (solomonzelman):

I'll give you a couple of examples of how graph is shifted for some different functions.

OpenStudy (solomonzelman):

Here, you go, `Example 1`. \(\large\color{ black }{\large {\bbox[5pt, white ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x \normalsize\color{red }{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x} \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ \underline{ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x} \normalsize\color{red}{ -~\rm{c} } }&~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (anonymous):

Thank you.. but what about number 1

OpenStudy (solomonzelman):

Here you go with another one, \(\large\color{ blue }{\large {\bbox[5pt, lightcyan ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ -~\rm{c} })^2 &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ +~\rm{c} })^2&~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ \underline{f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ -~\rm{c} }} &~\underline{ \rm{down~} } \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

which one? can you remind me please?

OpenStudy (anonymous):

If f(x) is the total cost, in dollars, of x candies, which of the following statements best describes the meaning of f(2) = 6? A. The total cost of 2 candies is $6. B. The total cost of 6 candies is $2. C. The total cost of 2 candies is $3. D. The total cost of 3 candies is $2.

OpenStudy (solomonzelman):

which one do you think? Do you have an opinion? Hint: `f(2) = 6` means that, when you plug in 2 for x, you get 6.

OpenStudy (anonymous):

The one i put was A

OpenStudy (solomonzelman):

Yes.

OpenStudy (anonymous):

Am i correct ?

OpenStudy (solomonzelman):

Yes you are:)

OpenStudy (anonymous):

Ok thank you a lot

OpenStudy (solomonzelman):

Anytime...

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of 3/2. A. y - 5 = 3/2(x + 1) B. y + 1 = 3/2(x - 5) C. y + 5 = 3/2(x - 1) D. y - 1 = 3/2(x + 5)